# Forecasting and Manipulating the Forecasts of Others

### 12 Mar 2026 v1

@@ A.3.2 Formal characterization
$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
@@ Appendix B Control First Order Conditions
This appendix derives the stationarity condition (B.8) and the closed backward system for the adjoint kernels $(\bar{H}^{X},H^{X},\{\bar{H}^{k},H^{k}\}_{k\neq i})$. By the innovations exogeneity established in Corollary 3.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic kernels weighting the exogenous innovation $I^{i}$. [^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [16, Theorem 7.12]; Lemma 3.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 3.5.
@@ B.1 The best-response problem in innovations coordinates
Fix a baseline deterministic-kernel Volterra profile and a deviating player $i$. By the innovations exogeneity established in Corollary 3.5, $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$, and $I^{i}$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion invariant to $D^{i}$. Player $i$’s control
$D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u)$
is a deterministic affine functional of $I^{i}$, and the optimization is over $(\bar{D}^{i},D^{i})$ with no self-referential signal dependence. Since the state dynamics are linear in $D^{i}$ and the cost is quadratic with $G^{i}_{DD}(t)\succ 0$, the induced objective is strictly convex: the stationarity condition derived below is the unique global optimum.
@@ B.2 Spike variation and the first-variation system
$\partial_{s}\delta w_{s}^{k}(u)=\widetilde{X}^{k}(s,u)^{\top}P^{k}(s)\Bigl(\delta X_{s}-\int_{0}^{s}X(s,r)\,\delta w_{s}^{k}(r)\,dr\Bigr),\qquad 0\leq u\leq s.$ (B.1)
The opponent control variation is $\delta D_{s}^{k}=\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du$, giving the coupled state system
$\frac{d}{ds}\delta X_{s}=A(s)\,\delta X_{s}+\sum_{k\neq i}\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du,\qquad\delta X_{t}=v.$ (B.2)
@@ B.3 Transition kernels
$\frac{\partial}{\partial t}\Phi^{XX}(s,t)$ $=-\Phi^{XX}(s,t)A(t)-\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}(t,z)^{\top}P^{k}(t)\,dz,$ (B.3)
$\frac{\partial}{\partial t}\Phi^{Xk}(s,t,u)$ $=-\Phi^{XX}(s,t)D^{k}(t,u)+\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}(t,z)^{\top}P^{k}(t)\,X(t,u)\,dz,$ (B.4)
@@ B.4 First variation of costs and the adjoint kernels
Fix player $i$ and suppress player superscripts on weight matrices. Using $\delta X_{s}=\Phi^{XX}(s,t)v$ and the Volterra form of $X_{s}$, the normalized cost variation is
$\delta J_{t}^{i}(v)=2\,\mathbb{E}\!\left[\left.v^{\top}\Bigl(G_{DD}^{i}(t)D_{t}^{i}+\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,d\widehat{W}_{t}^{i}(u)\Bigr)\ \right|\mathcal{F}_{t}^{i}\right],$ (B.5)
@@
$\bar{H}^{X}(t)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G_{XX}(s)\bar{X}(s)+G_{X}(s)\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr],$ (B.6)
$H^{X}(t,u)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}G_{XX}(s)\,X(s,u)\,ds+\Phi^{XX}(T,t)^{\top}G^{i}_{XX}(T)\,X(T,u).$ (B.7)
@@
$G_{DD}^{i}(t)D_{t}^{i}+\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,d\widehat{W}_{t}^{i}(u)=0\qquad\text{a.s.,}$ (B.8)
with equilibrium coefficients $\bar{D}^{i}(t)=-G_{DD}^{i}(t)^{-1}\bar{H}^{X}(t)$ and $D^{i}(t,u)=-G_{DD}^{i}(t)^{-1}H^{X}(t,u)$.
Equivalently, $D_{t}^{i}=-(G^{i}_{DD}(t))^{-1}\mathbb{E}[\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,dW_{u}\mid\mathcal{F}_{t}^{i}]$: the control is the conditional expectation of the full-information action (a separation property).
@@ B.5 Closed backward system for the adjoint kernels
$\bar{H}^{k}(t,u)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G_{XX}(s)\bar{X}(s)+G_{X}(s)\bigr]\,ds+\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr],$ (B.9)
$H^{k}(t,u,r)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}G_{XX}(s)\,X(s,r)\,ds+\Phi^{Xk}(T,t,u)^{\top}G^{i}_{XX}(T)\,X(T,r).$ (B.10)
@@
$\frac{d}{dt}\bar{H}^{X}(t)$ $=-\bigl[G^{i}_{XX}(t)\bar{X}(t)+G^{i}_{X}(t)\bigr]-A(t)^{\top}\bar{H}^{X}(t)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k}(t,z)\,dz,$ (B.11)
$\frac{d}{dt}\bar{H}^{k}(t,u)$ $=-D^{k}(t,u)^{\top}\bar{H}^{X}(t)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k}(t,z)\,dz,$ (B.12)
$\frac{d}{dt}H^{X}(t,r)$ $=-G^{i}_{XX}(t)X(t,r)-A(t)^{\top}H^{X}(t,r)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,H^{k}(t,z,r)\,dz,$ (B.13)
$\frac{d}{dt}H^{k}(t,u,r)$ $=-D^{k}(t,u)^{\top}H^{X}(t,r)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,H^{k}(t,z,r)\,dz,$ (B.14)
with terminal conditions $\bar{H}^{X}(T)=G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)$,  $\bar{H}^{k}(T,\cdot)=0$,  $H^{X}(T,r)=G^{i}_{XX}(T)X(T,r)$,  $H^{k}(T,\cdot,r)=0$.
@@ B.6 Risk-sensitive extension: exponential utility FOC
Under a deterministic-kernel Volterra profile, the best-response FOC is identical to Corollary 3.5 with $\widehat{W}_{t}^{i}$ replaced by the risk-adjusted noise-state
@@
where $C_{t}:=\mathsf{C}_{t}^{i}(u,s)=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional covariance kernel of $W$ given $\mathcal{F}_{t}^{i}$ (deterministic by Gaussianity of the conditional law under the Volterra profile), $K_{t}$ is the symmetric quadratic kernel of $\mathcal{C}^{i}$ in the Wiener-chaos representation (B.17), and $k_{t}$ is the corresponding linear kernel (B.16). The formula is valid when $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$, which holds for $\theta_{i}$ sufficiently small or $T$ sufficiently short. The filtering structure is unchanged; separation fails because the quadratic cost kernel rescales conditional precision via $(I-\theta_{i}C_{t}K_{t})^{-1}$ while $k_{t}$ shifts the conditional mean.
@@
Expanding $\mathcal{C}^{i}$ via the Volterra forms $X_{s}=\bar{X}(s)+\int_{0}^{s}X(s,u)\,dW_{u}$ and $D^{i}_{s}=\bar{D}^{i}(s)+\int_{0}^{s}\mathcal{D}^{i}(s,u)\,dW_{u}$ and applying stochastic Fubini yields the Wiener-chaos representation
@@
$\ell(r)$ $:=\int_{r}^{T}\!\Big[X(t,r)^{\top}\bigl(G^{i}_{XX}(t)\bar{X}(t)+G^{i}_{X}(t)\bigr)+\mathcal{D}^{i}(t,r)^{\top}G^{i}_{DD}(t)\,\bar{D}^{i}(t)\Big]\,dt$
$\qquad+\;X(T,r)^{\top}\bigl(G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr),$ (B.16)
$K(r,s)$ $:=\int_{\max(r,s)}^{T}\!\Big[X(t,r)^{\top}G^{i}_{XX}(t)\,X(t,s)+\mathcal{D}^{i}(t,r)^{\top}G^{i}_{DD}(t)\,\mathcal{D}^{i}(t,s)\Big]\,dt+X(T,r)^{\top}G^{i}_{XX}(T)\,X(T,s).$ (B.17)
@@
$G^{i}_{DD}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}\!\left[\mathcal{M}^{i}_{t}\,\Big|\,\mathcal{F}_{t}^{i}\right]=0,\qquad\mathbb{E}^{Q^{i}}[\,\cdot\mid\mathcal{F}_{t}^{i}]:=\frac{\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]}{\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]},$
@@
provided $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$ (satisfied for $|\theta_{i}|$ small or $T$ short). Rearranging gives
@@
$\boxed{G^{i}_{DD}(t)D^{i}_{t}\;+\;\bar{H}^{X}(t)\;+\;\int_{0}^{T}H^{X}(t,r)\,d_{r}\,\widehat{W}^{\,i,\theta}_{t}(r)\;=\;0,}$ (B.18)
@@
The first variation $\delta J^{i}_{\theta_{i},t}(v)$ is computed by differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude (Section B.2), which introduces the exponential tilt $Z^{i}:=e^{\theta_{i}\mathcal{C}^{i}}/\mathbb{E}[e^{\theta_{i}\mathcal{C}^{i}}]$. Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{i}_{DD}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure. Since $\mathcal{C}^{i}$ is quadratic in $W$ under the Volterra profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
@@
$M^{i}_{t}:=\bar{H}^{i}_{X}(t)+\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}W_{u},$
where $(\bar{H}^{i}_{X},H^{i}_{X})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
$D^{i,\mathrm{full}}_{t}$ $=-(G^{i}_{DD}(t))^{-1}M^{i}_{t},$ (B.20)
$D^{i}_{t}$ $=-(G^{i}_{DD}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
@@
$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
@@
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{i}_{DD}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
To verify that only $G^{i}_{DD}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$. The zero-order term vanishes by definition. All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{i}_{X})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$. The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{i}_{DD}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal). The standard LQG completion-of-the-square identifies these: using $G^{i}_{DD}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).

### 18 Mar 2026 v2

@@ A.3.2 Formal characterization
$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
@@ Appendix B Control First Order Conditions
This appendix derives the stationarity condition (B.8) and the closed backward system for the adjoint kernels $(\bar{H}^{X},H^{X},\{\bar{H}^{k},H^{k}\}_{k\neq i})$. By the innovations exogeneity established in Corollary 3.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic impulse-response maps weighting the exogenous innovation $I^{i}$. [^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [16, Theorem 7.12]; Lemma 3.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 3.5.
@@ B.1 The best-response problem in innovations coordinates
Fix a baseline noise-state linear profile and a deviating player $i$. By the innovations exogeneity established in Corollary 3.5, $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$, and $I^{i}$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion invariant to $D^{i}$. Player $i$’s control
$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$
is a deterministic affine functional of $I^{i}$, and the optimization is over $(\bar{D}^{i},D^{i})$ with no self-referential signal dependence. Since the state dynamics are linear in $D^{i}$ and the cost is quadratic with $G^{DD,i}(t)\succ 0$, the induced objective is strictly convex: the stationarity condition derived below is the unique global optimum.
@@ B.2 Spike variation and the first-variation system
$\partial_{s}\delta w_{s}^{k}(u)=\widetilde{X}^{k}_{s}(u)^{\top}P^{k}(s)\Bigl(\delta X_{s}-\int_{0}^{s}X_{s}(r)\,\delta w_{s}^{k}(r)\,dr\Bigr),\qquad 0\leq u\leq s.$ (B.1)
The opponent control variation is $\delta D_{s}^{k}=\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du$, giving the coupled state system
$\frac{d}{ds}\delta X_{s}=A(s)\,\delta X_{s}+\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,\qquad\delta X_{t}=v.$ (B.2)
@@ B.3 Transition kernels
$\frac{\partial}{\partial t}\Phi^{XX}(s,t)$ $=-\Phi^{XX}(s,t)A(t)-\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}_{t}(z)^{\top}P^{k}(t)\,dz,$ (B.3)
$\frac{\partial}{\partial t}\Phi^{Xk}(s,t,u)$ $=-\Phi^{XX}(s,t)D^{k}_{t}(u)+\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}_{t}(z)^{\top}P^{k}(t)\,X_{t}(u)\,dz,$ (B.4)
@@ B.4 First variation of costs and the adjoint kernels
Fix player $i$ and suppress player superscripts on weight matrices. Using $\delta X_{s}=\Phi^{XX}(s,t)v$ and the linear form of $X_{s}$, the normalized cost variation is
$\delta J_{t}^{i}(v)=2\,\mathbb{E}\!\left[\left.v^{\top}\Bigl(G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,d\widehat{W}_{t}^{i}(u)\Bigr)\ \right|\mathcal{F}_{t}^{i}\right],$ (B.5)
@@
$\bar{H}^{X}_{t}$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G^{XX}(s)\bar{X}_{s}+\bar{G}^{X,i}_{s}\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr],$ (B.6)
$H^{X}_{t}(u)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G^{XX}(s)\,X_{s}(u)+G^{X,i}_{s}(u)\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{XX,i}(T)\,X_{T}(u)+G^{X,i}_{T}(u)\bigr].$ (B.7)
@@
$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0\qquad\text{a.s.,}$ (B.8)
with equilibrium coefficients $\bar{D}^{i}_{t}=-G^{DD,i}(t)^{-1}\bar{H}^{X}_{t}$ and $D^{i}_{t}(u)=-G^{DD,i}(t)^{-1}H^{X}_{t}(u)$.
Equivalently, $D_{t}^{i}=-(G^{DD,i}(t))^{-1}\mathbb{E}[\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,dW_{u}\mid\mathcal{F}_{t}^{i}]$: the control is the conditional expectation of the full-information action (a certainty-equivalence property).
@@ B.5 Closed backward system for the adjoint kernels
$\bar{H}^{k}_{t}(u)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G^{XX}(s)\bar{X}_{s}+\bar{G}^{X,i}_{s}\bigr]\,ds+\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr],$ (B.9)
$H^{k}_{t}(u,r)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G^{XX}(s)\,X_{s}(r)+G^{X,i}_{s}(r)\bigr]\,ds$
$\qquad+\;\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{XX,i}(T)\,X_{T}(r)+G^{X,i}_{T}(r)\bigr].$ (B.10)
@@
$\frac{d}{dt}\bar{H}^{X}_{t}$ $=-\bigl[G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\bigr]-A(t)^{\top}\bar{H}^{X}_{t}-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k}_{t}(z)\,dz,$ (B.11)
$\frac{d}{dt}\bar{H}^{k}_{t}(u)$ $=-D^{k}_{t}(u)^{\top}\bar{H}^{X}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k}_{t}(z)\,dz,$ (B.12)
$\frac{d}{dt}H^{X}_{t}(r)$ $=-\bigl[G^{XX,i}(t)X_{t}(r)+G^{X,i}_{t}(r)\bigr]-A(t)^{\top}H^{X}_{t}(r)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,H^{k}_{t}(z,r)\,dz,$ (B.13)
$\frac{d}{dt}H^{k}_{t}(u,r)$ $=-D^{k}_{t}(u)^{\top}H^{X}_{t}(r)+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,H^{k}_{t}(z,r)\,dz,$ (B.14)
with terminal conditions $\bar{H}^{X}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$,  $\bar{H}^{k}_{T}(\cdot)=0$,  $H^{X}_{T}(r)=G^{XX,i}(T)X_{T}(r)+G^{X,i}_{T}(r)$,  $H^{k}_{T}(\cdot,r)=0$.
@@ B.6 Risk-sensitive extension: exponential utility FOC
Under a linear profile, the best-response FOC is identical to Corollary 3.5 with $\widehat{W}_{t}^{i}$ replaced by the risk-adjusted noise-state
@@
where $C_{t}:=\mathsf{C}_{t}^{i}(u,s)=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional covariance kernel of $W$ given $\mathcal{F}_{t}^{i}$ (deterministic by Gaussianity of the conditional law under the linear profile), $K_{t}$ is the symmetric quadratic kernel of $\mathcal{C}^{i}$ in the Wiener-chaos representation (B.17), and $k_{t}$ is the corresponding linear kernel (B.16). The formula is valid when $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$, which holds for $\theta_{i}$ sufficiently small or $T$ sufficiently short. The filtering structure is unchanged; certainty equivalence fails because the quadratic cost kernel rescales conditional precision via $(I-\theta_{i}C_{t}K_{t})^{-1}$ while $k_{t}$ shifts the conditional mean.
@@
Expanding $\mathcal{C}^{i}$ via the linear forms $X_{s}=\bar{X}_{s}+\int_{0}^{s}X_{s}(u)\,dW_{u}$ and $D^{i}_{s}=\bar{D}^{i}_{s}+\int_{0}^{s}\mathcal{D}^{i}_{s}(u)\,dW_{u}$ and applying stochastic Fubini yields the Wiener-chaos representation
@@
$\ell(r)$ $:=\int_{r}^{T}\!\Big[X_{t}(r)^{\top}\bigl(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\bigr)+G^{X,i}_{t}(r)^{\top}\bar{X}_{t}$
$\qquad\qquad+\mathcal{D}^{i}_{t}(r)^{\top}G^{DD,i}(t)\,\bar{D}^{i}_{t}\Big]\,dt$
$\qquad+\;X_{T}(r)^{\top}\bigl(G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr)+G^{X,i}_{T}(r)^{\top}\bar{X}_{T},$ (B.16)
$K(r,s)$ $:=\int_{\max(r,s)}^{T}\!\Big[X_{t}(r)^{\top}G^{XX,i}(t)\,X_{t}(s)+X_{t}(r)^{\top}G^{X,i}_{t}(s)$
$\qquad\qquad+G^{X,i}_{t}(r)^{\top}X_{t}(s)+\mathcal{D}^{i}_{t}(r)^{\top}G^{DD,i}(t)\,\mathcal{D}^{i}_{t}(s)\Big]\,dt$
$\qquad+\;X_{T}(r)^{\top}G^{XX,i}(T)\,X_{T}(s)+X_{T}(r)^{\top}G^{X,i}_{T}(s)+G^{X,i}_{T}(r)^{\top}X_{T}(s).$ (B.17)
@@
$G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}\!\left[\mathcal{M}^{i}_{t}\,\Big|\,\mathcal{F}_{t}^{i}\right]=0,\qquad\mathbb{E}^{Q^{i}}[\,\cdot\mid\mathcal{F}_{t}^{i}]:=\frac{\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]}{\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]},$
@@
provided $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$ (satisfied for $|\theta_{i}|$ small or $T$ short). Here $C_{t}$ and $K_{t}$ act as integral operators on $L^{2}([0,t];\mathbb{R}^{(n+1)d})$, so $(I-\theta_{i}C_{t}K_{t})^{-1}$ is the resolvent (Neumann series) $(I-\theta_{i}C_{t}K_{t})^{-1}f=f+\theta_{i}C_{t}K_{t}f+\theta_{i}^{2}(C_{t}K_{t})^{2}f+\cdots$, convergent precisely when $\|\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\|_{\mathrm{op}}<1$. Rearranging gives
@@
$\boxed{G^{DD,i}(t)D^{i}_{t}\;+\;\bar{H}^{X}_{t}\;+\;\int_{0}^{T}H^{X}_{t}(r)\,d_{r}\,\widehat{W}^{\,i,\theta}_{t}(r)\;=\;0,}$ (B.18)
@@
The first variation $\delta J^{i}_{\theta_{i},t}(v)$ is computed by differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude (Section B.2), which introduces the exponential tilt $Z^{i}:=e^{\theta_{i}\mathcal{C}^{i}}/\mathbb{E}[e^{\theta_{i}\mathcal{C}^{i}}]$. Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure. Since $\mathcal{C}^{i}$ is quadratic in $W$ under the linear profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
@@
$M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u},$
where $(\bar{H}^{X,i},H^{X,i})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
$D^{i,\mathrm{full}}_{t}$ $=-(G^{DD,i}(t))^{-1}M^{i}_{t},$ (B.20)
$D^{i}_{t}$ $=-(G^{DD,i}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
@@
$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
@@
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{DD,i}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
To verify that only $G^{DD,i}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$. The zero-order term vanishes by definition. All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{X,i})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$. The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{DD,i}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal). The standard LQG completion-of-the-square identifies these: using $G^{DD,i}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).